OpenStudy (sette):

Dont Understandd . .. Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2.

OpenStudy (sette):

it starts out :Y= x^2+4x-

OpenStudy (mertsj):

y=a(x-h)^2+k (h,k) is the vertex and for x and y we can use one of the x intercepts. Let's choose (2,0) 0=a(2+2)^2-16 0=16a-16 a=1

OpenStudy (mertsj):

y=1(x+2)^2-16 is the required equation. If you want it in a different form, it can be rearranged.

OpenStudy (mertsj):

y=x^2+4x+4-16 y=x^2+4x-12

OpenStudy (sette):

That makes sense. Thank You .